I tried to do this three times, but get lost in all the steps. Can you help me?
C(10,5)x C(15,8)
________________
C(12,11)

Answer by checkley75(3666)

C(10,5)x C(15,8)
________________
C(12,11)
10C5=252
15C8=5,435
12C11=12
252*6,435/12=1,621,620/12=135,135 ANSWER.